OpenStudy (anonymous):

Find the difference: (2x² − 5x + 2) − (x² − 2x + 5)

OpenStudy (theviper):

\[(2x² − 5x + 2) − (x² − 2x + 5)\]\[=2x² − 5x + 2-x^2+2x-5\]\[=x^2-3x-3\]

OpenStudy (theviper):

Understood??

OpenStudy (anonymous):

yep

OpenStudy (theviper):

nice :)